# Trained Random Forests Completely Reveal your Dataset

### v1

@@ 1 Introduction
Machine learning (ML) techniques are increasingly used on sensitive data, such as medical records for kidney exchange (Aziz et al., 2021), criminal records (Angwin et al., 2016) or credit history. As this raises significant ethical and societal challenges, the use of such private data is directly regulated by several legal texts, such as the recent European Union General Data Protection Regulation [^1] https://gdpr-info.eu/ or the forthcoming AI Act [^7] https://artificialintelligenceact.eu/. Privacy has attracted significant attention during the last decades (Liu et al., 2021a) in order to protect sensitive or personal information about individual users while still being able to extract useful patterns from data. Moreover, privacy risks may further be exacerbated by the consideration of other ethical desiderata, e.g., while releasing a trained ML model for the sake of transparency.
@@ 2 Technical Background
The training dataset is used to build a random forest $\mathcal{T}$ in which each tree $t\in\mathcal{T}$ is made of a set of internal nodes $\mathcal{V}_{t}^{I}$ and a set of leaves $\mathcal{V}_{t}^{L}$. Each internal node $v\in\mathcal{V}_{t}^{I}$ corresponds to a binary condition over the value of a given attribute. If the condition is satisfied, the example being classified descends towards the left child $l(v)$ of the node, otherwise it descends towards its right child $r(v)$. Once the example reaches a leaf $v\in\mathcal{V}_{t}^{L}$ (terminal node), it is classified according to the class associated with this leaf. Such class corresponds to the majority class among the training examples captured by the leaf. To compute it (and eventually assign class probabilities), each leaf contains the per-class number of training examples it captures. In popular ML libraries such as scikit-learn, such counts are also provided in the internal nodes, as shown in Figure 1. Then, for every node $v\in\mathcal{V}_{t}^{I}\bigcup\mathcal{V}_{t}^{L}$, let $nb_{tvc}$ denote the number of training examples of class $c$ that went through $v$.
@@ 3 Related Works
ML methods often exploit private data during training. Consequently, it is crucial to ensure that their outputs – which may be released directly or accessed through a dedicated API – do not leak information regarding their inputs (Dinur & Nissim, 2003). Inference attacks against ML models (Rigaki & Garcia, 2020) precisely aim at exploiting the output of a learning algorithm to infer information regarding the training dataset. Different attacks can be distinguished, depending on their specific objective. For instance, membership inference attacks aim to infer whether an example was part of a model’s training data or not (Shokri et al., 2017; Carlini et al., 2022). In this work, we are interested in dataset reconstruction attacks, which aim at reconstructing (entirely or partially) a model’s training dataset (Dwork et al., 2017). We consider the white-box setup, in which the adversary has complete knowledge of the model’s parameters instead of a black-box API access to it (Cristofaro, 2020). Moreover, we specifically target RF models.
@@
Reconstruction attacks are one of the most ambitious inference attacks against ML models, as they directly aim to recover entire parts of the training data. However, instead of attempting to reconstruct the whole training set, most reconstruction attacks only target retrieving part of it. For instance, the first reconstruction attacks (originally proposed against database access mechanisms) only aimed at retrieving one private binary attribute for all the database examples - assuming all other attributes were publicly known (Dinur & Nissim, 2003; Dwork et al., 2007, 2017). Some other studies only target reconstructing part of one particular example, given some public information about it (Fredrikson et al., 2014, 2015). Other approaches require additional knowledge, such as intermediate gradients computed during collaborative (Phong et al., 2017) or online (Salem et al., 2020) learning, or information related to the model’s fairness (Hu & Lan, 2020; Aalmoes et al., 2022; Hamman et al., 2022; Ferry et al., 2023).
@@ 4 Illustrative Example
We first give an intuition of the reconstruction problem on a small dataset (Table 1) with $4$ examples described by $4$ binary attributes $f_{i\in\{1..4\}}$ and a binary class $c$. Figure 1 provides two decision trees trained on this dataset. Tree $t_{1}$ was trained without using $f_{1}$, while tree $t_{2}$ was trained without using $f_{3}$ (though this information is unknown to the reconstruction algorithm). For presentation simplicity, bagging is not used here, and therefore each training example is used a single time in each tree. By following the paths from the root to each leaf within $t_{1}$ (Figure 0(a)), one can set the value of some attributes within the reconstructed dataset by leveraging the performed splits and the per-node cardinalities. For instance, following the leftmost path, we can observe that the two examples belonging to class 00 have value 00 for both $f_{3}$ and $f_{2}$. Such information permits to directly fix some attributes’ values. In a similar fashion, according to $t_{2}$, there exists exactly one example of class $1$ with value 00 for $f_{2}$, and another one with value $1$.
@@
The main issue with such manual reasoning is that, except in some obvious cases (i.e., when all examples of a certain class respect a given splitting condition), splits will permit quantifying how many examples respect a certain condition without telling which are these examples. Therefore, the biggest challenge of dataset reconstruction is to individually link the examples between the different trees and find a compatible dataset that respects all the cardinality constraints. This challenge is exacerbated by the bagging process, as in this case, the cardinalities within the trees’ nodes may count some examples several times (and ignore some others).
@@ 5 NP-Hardness Result
In this section, we formally define the dataset reconstruction problem (DRP) and show its $\mathcal{NP}$-completeness. Part of the input data for (DRP) has already been defined before: the set of classes $\mathcal{C}$, the number $N$ of examples, the number $M$ of binary attributes, and the forest $\mathcal{T}$, where each $t\in\mathcal{T}$ is a binary tree. Also, for each class $c\in\mathcal{C}$, tree $t\in\mathcal{T}$ and each $v\in\mathcal{V}_{t}^{I}\cup\mathcal{V}_{t}^{L}$, we are given an amount $nb_{tvc}\in\mathbb{Z}_{+}$ of examples of class $c$ that are classified in node $v$ of tree $t$.
@@
  • lower/upper bounds $l_{k},u_{k}\in\mathbb{Z}_{+},\forall k\in\{1..N\}$.
@@
  • For each class $c\in\mathcal{C}$, tree $t\in\mathcal{T}$ and each $v\in\mathcal{V}_{t}^{I}$, we have that $nb_{tvc}=nb_{tl(v)c}+nb_{tr(v)c}$.
  • For each tree $t\in\mathcal{T}$, we have $\sum\limits_{v\in\mathcal{V}_{t}^{L}}\sum\limits_{c\in\mathcal{C}}nb_{tvc}=N$.
For each tree $t\in\mathcal{T}$, we let $\Phi_{v}\subseteq\{1..M\}$ denote the set of indices of the boolean attributes that must be True for an example to fall into node $v\in\mathcal{V}_{t}^{L}\cup\mathcal{V}_{t}^{I}$. Similarly, $\overline{\Phi_{v}}\subseteq\{1..M\}$ is the set of indices of the boolean attributes that must be False for an example to fall into $v$ (hence $\Phi_{v}\cap\overline{\Phi_{v}}=\emptyset$). Both represent the splits that are found along the path from the root node of tree $t\in\mathcal{T}$ to $v$. Formally, if $v$ is the root node of the tree, then $\Phi_{v}=\overline{\Phi_{v}}=\emptyset$. For every $v\in\mathcal{V}_{t}^{I}$, we can define such sets for its children as $\Phi_{l(v)}=\Phi_{v}$ and $\overline{\Phi_{{l(v)}}}=\overline{\Phi_{v}}\cup\{f_{v}\}$ ; $\Phi_{r(v)}=\Phi_{v}\cup\{f_{v}\}$ and $\overline{\Phi_{{r(v)}}}=\overline{\Phi_{v}}$
@@
  • $\sum\limits_{v\in\mathcal{V}_{t}^{L}}y_{tvk}\in[l_{k},u_{k}]$, $\forall t\in\mathcal{T},k\in\{1..N\}$
  • $\sum\limits_{k\in\{1,\ldots,N\}:z_{k}=c}y_{tvk}=nb_{tvc}$, $\forall t\in\mathcal{T},v\in\mathcal{V}_{t}^{L}$
  • For all $k\in\{1..N\}$, $t\in\mathcal{T}$, $v\in\mathcal{V}_{t}^{L}$, if $y_{tvk}>0$, then $(x_{k})_{i}=0$ for all $i\in\overline{\Phi_{v}}$ and $(x_{k})_{i}=1$ for all $i\in\Phi_{v}$.
We note that the last constraint implies that for every $t\in\mathcal{T}$ and $k\in\{1..N\}$, at most one variable in the set $\{y_{tvk}:v\in\mathcal{V}_{t}^{L}\}$ is nonzero. This is due to the fact that, from the way $\Phi_{v}$ and $\overline{\Phi_{v}}$ are constructed, all other leaves $v^{\prime}\in\mathcal{V}_{t}^{L}\setminus\{v\}$ must have at least one attribute $i\in\Phi_{v}\cup\overline{\Phi_{v}}$ switching its required True/False value in $v^{\prime}$.
Note that if we set $l_{k}=u_{k}=1,\forall k\in\{1..N\}$, we impose that each example must appear exactly once in every tree, which corresponds to the situation when no bagging is used. If we set $l_{k}=0,u_{k}=N,\forall k\in\{1..N\}$, we get that each example can appear any number of times in a tree, which corresponds to the situation when bagging is used.
@@
(DRP) is $\mathcal{NP}$-complete, even for $N=1$, $|\mathcal{C}|=1$ and trees of depth at most 3.
@@
First, the feasibility of any solution $(x,z,y)$ can be checked in polynomial time, so $(DRP)\in\mathcal{NP}$.
Next, consider an instance of the $\mathcal{NP}$-complete problem 3-SAT, given by a set of $\mathcal{T}$ clauses with $3$ literals each. Furthermore, each literal is one of $M$ possible ones. We construct an instance of (DRP) with $N=1$, $|\mathcal{C}|=1$ that can solve 3-SAT. For each clause $t\in\mathcal{T}$, build a perfect binary tree $t$ of depth 3. For every $v\in\mathcal{V}_{t}^{I}$ at depth $d$, let $f_{v}$ be the $(d+1)$-th literal of clause $t$.
In this way, the fixed attributes $\Phi_{v}$ and $\overline{\Phi_{v}}$ of the leaves $v$ of the tree $t$ will correspond to every possible assignment of the literals in $t$ to 0 or 1. So we set $nb_{tvc}=1$ for the leaf corresponding to the assignment of literals that would make clause $t$ true, and 0 to all other leaves. The other $nb_{tvc}$ can be set to satisfy the assumptions on the data.
With this, we observe that (DRP) is feasible if and only if one can assign literals $x_{11},..,x_{1M}$ to satisfy all clauses in $\mathcal{T}$. Note that it does not matter what values $l_{k}$ and $u_{k}$ are set to, as long as $l_{k}\leq 1\leq u_{k}$, so this proves hardness for both models with and without bagging. ∎
@@ 6 Constraint Programming Approach
As seen in Section 4, an inspection of the different trees gives sets of restrictions over feature values that concern a known number of examples of each class. However, it does not tell which example specifically satisfies which condition. Testing feasible combinations by inspection would require extensive trial and error, leading to an intractable process. Instead, we propose to formulate this search problem as a constraint programming (CP) model, permitting the use of efficient out-of-the-shelf solvers for such models. The model we design covers the most general case where bagging is used to train the forest, and includes discretization strategies specifically designed to help the solution process. Note that, while we focus on CP, Mixed-Integer Linear Programming (MILP) could also be employed instead. However, having conducted experiments with both techniques, and as demonstrated in Appendix A, CP generally achieved better performance and permitted to handle bagging much more effectively.
@@
To define the objective function, we will assume that a training example can only appear at most $7$ times in any tree [^1] Moreover, if no feasible solution is found using such value, it is still possible to increase it and run the solver again.. With the default parameters of scikit-learn, the probability of any example appearing more than $7$ times in a tree is less than $10^{-4}$. With this, we define :=$\mathcal{B}:=\{0,\ldots,7\}$. We now define a binary variable to capture how many times an example is used:
@@
  • $\forall t\in\mathcal{T},\forall v\in\mathcal{V}_{t}^{L},\forall c\in\mathcal{C}:nb_{tvc}=\sum\limits_{{k\in\{1..N\}}}y_{tvkc}$
@@
    if $\textbf{if }\sum\limits_{c\in\mathcal{C}}y_{tvkc}\geq 1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\overline{\Phi_{v}}}{x_{ki}=0}\right)$
@@
RFs can be trained using random subsets of features for each split but considering all the examples in each tree. In such situations without bagging, the CP model can be significantly simplified. Variables $y_{tvkc}$ will become binary and sum up to $1$ for each tree $t\in\mathcal{T}$ and each example $k\in\{1..N\}$, since each example will be used exactly once in each tree. Also, we know in advance how many examples are of each class $c\in\mathcal{C}$. Therefore, to match that data, we may fix the variables $z_{kc}$ in advance. Finally, $q_{tkb}$ are always fixed since every example is used exactly once in each tree. Thus, the objective function becomes constant and the problem reduces to the search of a feasible solution.

### v2

@@ 1 Introduction
Machine learning (ML) techniques are increasingly used on sensitive data, such as medical records for kidney exchange (Aziz et al., 2021), criminal records (Angwin et al., 2016) or credit history. As this raises significant ethical and societal challenges, the use of such private data is directly regulated by several legal texts, such as the recent European Union General Data Protection Regulation [^1] https://gdpr-info.eu/ or the forthcoming AI Act [^7] https://artificialintelligenceact.eu/. Privacy has attracted significant attention during the last decades (Liu et al., 2021a) in order to protect sensitive or personal information about individual users while still being able to extract useful patterns from data. Moreover, privacy risks may further be exacerbated by the consideration of other ethical desiderata, e.g., when releasing a trained ML model for the sake of transparency.
@@ 2 Technical Background
The training dataset is used to build a random forest $\mathcal{T}$ in which each tree $t\in\mathcal{T}$ is made of a set of internal nodes $\mathcal{V}_{t}^{I}$ and a set of leaves $\mathcal{V}_{t}^{L}$. Each internal node $v\in\mathcal{V}_{t}^{I}$ corresponds to a binary condition over the value of a given attribute. If the condition is satisfied, the example being classified descends towards the left child $l(v)$ of the node, otherwise it descends towards its right child $r(v)$. Once the example reaches a leaf $v\in\mathcal{V}_{t}^{L}$ (terminal node), it is classified according to the class associated with this leaf. Such class corresponds to the majority class among the training examples captured by the leaf. To compute it (and eventually assign class probabilities), each leaf contains the per-class number of training examples it captures. In popular ML libraries such as scikit-learn, such counts are also provided in the internal nodes, as shown in Figure 1. Then, for every node $v\in\mathcal{V}_{t}^{I}\bigcup\mathcal{V}_{t}^{L}$, let $n_{tvc}$ denote the number of training examples of class $c$ that went through $v$.
@@ 3 Related Works
ML methods often exploit private data during training. Consequently, it is crucial to ensure that their outputs – which may be released directly or accessed through a dedicated API – do not leak information regarding their inputs (Dinur & Nissim, 2003). Inference attacks against ML models (Rigaki & Garcia, 2020) precisely aim at exploiting the output of a learning algorithm to infer information regarding the training dataset. Different attacks can be distinguished, depending on their specific objective. For instance, membership inference attacks aim to infer whether an example was part of a model’s training data or not (Shokri et al., 2017; Carlini et al., 2022). In this work, we are interested in dataset reconstruction attacks, which aim at reconstructing (entirely or partially) a model’s training dataset (Dwork et al., 2017). The considered attack model is as follows. We specifically target RF models, and consider the white-box setup, in which the adversary has complete knowledge of the model’s parameters instead of a black-box API access to it (Cristofaro, 2020). He also knows the domains of the different attributes involved in the data. Importantly, he does not intervene during the training process, but rather gets the trained model afterwards.
@@
Reconstruction attacks are one of the most ambitious inference attacks against ML models, as they directly aim to recover entire parts of the training data. However, instead of attempting to reconstruct the whole training set, most reconstruction attacks only target retrieving part of it. For instance, the first reconstruction attacks (originally proposed against database access mechanisms) only aimed at retrieving one private binary attribute for all the database examples - assuming all other attributes were publicly known (Dinur & Nissim, 2003; Dwork et al., 2007, 2017). Some other studies only target reconstructing part of one particular example, given some public information about it (Fredrikson et al., 2014, 2015). Other approaches require additional knowledge, such as intermediate gradients computed during collaborative (Phong et al., 2017) or online (Salem et al., 2020) learning, stationary points reached by gradient descent algorithms (Haim et al., 2022) or information regarding the model’s fairness (Hu & Lan, 2020; Aalmoes et al., 2022; Hamman et al., 2022; Ferry et al., 2023).
@@ 4 Illustrative Example
We first give an intuition of the reconstruction problem on a small dataset (Table 1) with $4$ examples described by $4$ binary attributes $f_{i\in\{1..4\}}$ and a binary class $c$. Figure 1 provides two decision trees trained on this dataset. Tree $t_{1}$ was trained without using $f_{1}$, while tree $t_{2}$ was trained without using $f_{3}$ (though this information is unknown to the reconstruction algorithm). For presentation simplicity, bagging is not used here, and therefore each training example is used a single time in each tree. By following the paths from the root to each leaf within $t_{1}$ (Figure 1(a)), one can set the value of some attributes within the reconstructed dataset by leveraging the performed splits and the per-node cardinalities. For instance, following the leftmost path, we can observe that the two examples belonging to class 00 have value 00 for both $f_{3}$ and $f_{2}$. Such information permits to fix some attributes’ values directly. Similarly, according to $t_{2}$, there exists exactly one example of class $1$ with value 00 for $f_{2}$, and another one with value $1$.
The main issue with such ad-hoc reasoning is that, except in some obvious cases (i.e., when all examples of a certain class respect a given splitting condition), splits will permit quantifying how many examples respect a certain condition without telling which are these examples. Therefore, the biggest challenge of dataset reconstruction is to individually link the examples between the different trees and find a compatible dataset that respects all the cardinality constraints. This challenge is exacerbated by the bagging process, as in this case, the cardinalities within the trees’ nodes may count some examples several times (and ignore some others).
@@ 5 NP-Hardness Result
In this section, we formally define the dataset reconstruction problem (DRP) and show its $\mathcal{NP}$-completeness. Part of the input data for (DRP) has already been defined before: the set of classes $\mathcal{C}$, the number $N$ of examples, the number $M$ of binary attributes, and the forest $\mathcal{T}$, where each $t\in\mathcal{T}$ is a binary tree. Also, for each class $c\in\mathcal{C}$, tree $t\in\mathcal{T}$ and each $v\in\mathcal{V}_{t}^{I}\cup\mathcal{V}_{t}^{L}$, we are given an amount $n_{tvc}\in\mathbb{Z}_{+}$ of examples of class $c$ that are classified in node $v$ of tree $t$.
@@
  • a set $\mathcal{B}$ of integer values, that represent how many times a sample may appear in a tree
@@
  • For each class $c\in\mathcal{C}$, tree $t\in\mathcal{T}$ and each $v\in\mathcal{V}_{t}^{I}$, we have that $n_{tvc}=n_{tl(v)c}+n_{tr(v)c}$.
  • For each tree $t\in\mathcal{T}$, we have $\sum\limits_{v\in\mathcal{V}_{t}^{L}}\sum\limits_{c\in\mathcal{C}}n_{tvc}=N$.
For each tree $t\in\mathcal{T}$, we let $\Phi_{v}^{+}\subseteq\{1..M\}$ denote the set of indices of the boolean attributes that must be True for an example to fall into node $v\in\mathcal{V}_{t}^{L}\cup\mathcal{V}_{t}^{I}$. Similarly, $\Phi_{v}^{-}\subseteq\{1..M\}$ is the set of indices of the boolean attributes that must be False for an example to fall into $v$ (hence $\Phi_{v}^{+}\cap\Phi_{v}^{-}=\emptyset$). Both represent the splits that are found along the path from the root node of tree $t\in\mathcal{T}$ to $v$. Formally, if $v$ is the root node of the tree, then $\Phi_{v}^{+}=\Phi_{v}^{-}=\emptyset$. For every $v\in\mathcal{V}_{t}^{I}$, we can define such sets for its children as $\Phi_{l(v)}^{+}=\Phi_{v}^{+}$ and $\Phi_{{l(v)}}^{-}=\Phi_{v}^{-}\cup\{f_{v}\}$ ; $\Phi_{r(v)}^{+}=\Phi_{v}^{+}\cup\{f_{v}\}$ and $\Phi_{{r(v)}}^{-}=\Phi_{v}^{-}$
@@
  • $\sum\limits_{v\in\mathcal{V}_{t}^{L}}y_{tvk}\in\mathcal{B}$, $\forall t\in\mathcal{T},k\in\{1..N\}$
  • $\sum\limits_{k\in\{1,\ldots,N\}:z_{k}=c}y_{tvk}=n_{tvc}$, $\forall t\in\mathcal{T},v\in\mathcal{V}_{t}^{L}$
  • For all $k\in\{1..N\}$, $t\in\mathcal{T}$, $v\in\mathcal{V}_{t}^{L}$, if $y_{tvk}>0$, then $(x_{k})_{i}=0$ for all $i\in\Phi_{v}^{-}$ and $(x_{k})_{i}=1$ for all $i\in\Phi_{v}^{+}$.
We note that the last constraint implies that for every $t\in\mathcal{T}$ and $k\in\{1..N\}$, at most one variable in the set $\{y_{tvk}:v\in\mathcal{V}_{t}^{L}\}$ is nonzero. This is due to the fact that, from the way $\Phi_{v}^{+}$ and $\Phi_{v}^{-}$ are constructed, all other leaves $v^{\prime}\in\mathcal{V}_{t}^{L}\setminus\{v\}$ must have at least one attribute $i\in\Phi_{v}^{+}\cup\Phi_{v}^{-}$ switching its required True/False value in $v^{\prime}$.
Note that if we set =$\mathcal{B}=\{1\}$, we impose that each example must appear exactly once in every tree, which corresponds to the situation when no bagging is used. If we set =$\mathcal{B}=\{0,\ldots,N\}$, we get that each example can appear any number of times in a tree, which corresponds to the situation when bagging is used.
@@
The decision version of (DRP) is $\mathcal{NP}$-complete.
@@
First, a YES certificate of (DRP) is any solution $(x,z,y)$. One can verify if it is feasible in polynomial time, so the decision version of $(DRP)\in\mathcal{NP}$.
Next, consider an instance of the $\mathcal{NP}$-complete problem 3-SAT, given by a set $L=\{1,\ldots,|L|\}$ of clauses with three literals each. Each literal is either a variable or its complement, and there are $V$ possible variables. We construct an instance of (DRP) with $|\mathcal{C}|=1$, by specifying the forest $\mathcal{T}$ and $n_{lvc}$ values. We let =$\mathcal{B}=\{1\}$. By assumption on the data, we only need to specify $n_{lvc}$ for $v\in\mathcal{V}_{t}^{L}$. Also, recall that the left branch of $v\in\mathcal{V}_{t}^{I}$ corresponds to setting $f_{v}$ to zero, and the right branch sets it to 1. The idea is to construct a perfect binary tree for each $l\in L$. This way, the fixed attributes $\Phi_{v}^{+}$ and $\Phi_{v}^{-}$ of each leaf $v$ of tree $l$ will represent one of each possible assignment of the three literals of $l$ to 0 or 1.
Out of the eight possible leaves of tree $l$, seven satisfy the corresponding clause, and one does not. We set $n_{lvc}=1$ for the leaves that satisfy clause $l$ and $n_{lvc}=0$ otherwise. The goal of our construction is to force the DRP solver to generate an example whose attribute values lead it towards one of the leaves with $n_{lvc}=1$ in each tree, i.e., a 3-SAT solution, whenever such a solution exists. To achieve this, we include six additional “dummy” examples for each tree to reach the remaining alternatives with $n_{lvc}=1$. Therefore, our constructed instance has $N=6|L|+1$ examples.
We rely on $M=V+|L|$ binary attributes. The first $V$ attributes will represent the 0/1 assignment of values to literals. The remaining $|L|$ attributes will determine whether one example is used or not in the leaves corresponding to the perfect binary tree of that clause. With that, we add to each tree $l=1,\ldots,|L|$, a root node where we branch on feature $V+l$. The right branch of the tree will contain the perfect binary tree described above. The left branch is a single leaf node with $n_{lvc}=6|L|-6$, designed to absorb all dummy examples not destined toward that tree.
We finally construct one extra auxiliary tree. All its left branches end up in leaves. Each node on the right branch at depth $d$ (including the root node) branches on attribute $V+d+1$. The left leaf at depth 1 has $n_{lvc}=6|L|-6$. The left leaf at depth 2 has $n_{lvc}=6$. All other left leaves have $n_{lvc}=0$. The rightmost leaf has $n_{lvc}=1$. With this, the rightmost leaf imposes that a single example reaches one leaf of every perfect tree (the desired 3-SAT solution). The other nodes of the tree just count the extra examples.
Appendix A proves that (DRP) is feasible if and only if the 3-SAT instance is a YES instance. It also contains illustrative examples for the construction. ∎
@@ 6 Constraint Programming Approach
As seen in Section 4, an inspection of the different trees gives sets of restrictions over feature values that concern a known number of examples of each class. However, it does not tell which example specifically satisfies which condition. Testing feasible combinations by inspection would require extensive trial and error, leading to an intractable process. Instead, we propose to formulate this search problem as a constraint programming (CP) model, permitting the use of efficient out-of-the-shelf solvers for such models. The model we design covers the most general case where bagging is used to train the forest, and includes discretization strategies specifically designed to help the solution process. Note that, while we focus on CP, Mixed-Integer Linear Programming (MILP) could also be employed instead. However, having conducted experiments with both techniques, and as demonstrated in Appendix B, CP generally achieved better performance and permitted to handle bagging much more effectively.
@@
To define the objective function, we will assume that a training example appears at most $7$ times in any tree, since higher values are very unlikely. Indeed, bootstrap sampling consists of sampling with replacement $N$ examples from a set of $N$ original examples. At each iteration of the bootstrap sampling process, each example has a probability of $\frac{1}{N}$ of being selected. The probability of an example being selected more than B times can then be computed as:
$1-\sum\limits_{b=0}^{B-1}\left(\left(\frac{1}{N}\right)^{b}\cdot\left(\frac{N-1}{N}\right)^{N-b}\cdot\binom{N}{b}\right).$
If $N=100$ as in our experiments, the probability of an example appearing more than $7$ times in a bootstrap sampled training set is roughly $10^{-5}$. This value remains similar for larger values of $N$ (e.g., around $10^{-5}$ for $N=10^{10}$).
With this, we define :=$\mathcal{B}:=\{0,\ldots,7\}$. Note that if no feasible solution is found using this default value, increasing it and solving the model again is possible. We now define a binary variable to capture how many times an example is used:
@@
  • $\forall t\in\mathcal{T},\forall v\in\mathcal{V}_{t}^{L},\forall c\in\mathcal{C}:n_{tvc}=\sum\limits_{{k\in\{1..N\}}}y_{tvkc}$
@@
    if $\textbf{if }\sum\limits_{c\in\mathcal{C}}y_{tvkc}\geq 1\textbf{ then}\quad\left(\bigwedge\limits_{i\in\Phi_{v}^{+}}{x_{ki}=1}\right)\land\left(\bigwedge\limits_{i\in\Phi_{v}^{-}}{x_{ki}=0}\right)$
@@
RFs can be trained using random subsets of features for each split but considering all the examples in each tree. In such situations without bagging, the CP model can be significantly simplified. Variables $y_{tvkc}$ will become binary and sum up to $1$ for each tree $t\in\mathcal{T}$ and each example $k\in\{1..N\}$, since each example will be used exactly once in each tree. Also, we know in advance how many examples are of each class $c\in\mathcal{C}$. Therefore, to match that data, we may fix the variables $z_{kc}$ in advance. Finally, $q_{tkb}$ are always fixed since every example is used exactly once in each tree. Thus, the objective function becomes constant, and the problem reduces to the search for a feasible solution.
Reconstructing non-binary attributes.
To streamline the presentation of the methodology and evaluation metric, we provided our model for the particular case of binary attributes. However, extending it to handle other types of attributes is possible with minimal changes, as detailed in Appendix D and implemented within our publicly available repository. In a nutshell, categorical attributes are typically one-hot encoded for tree ensembles and directly handled by our formulation. Ordinal features can be modeled as integer variables and only require a slight generalization of the constraints connecting the attributes’ values to the assignment of the examples to the leaves. Finally, numerical attributes can also be reconstructed: although they take values in a continuous space, the number of splits within the forest is finite, and so is the number of different intervals in which they can lie. Leveraging this observation, we can use ordinal features to model such possible intervals in the reconstruction.